# Absolute abstraction: a renormalisation group approach

### v4

@@ 5 Discussion
The main original contribution of this paper is to argue that universal representations emerge spontaneously from the combined effects of depth and breadth. We derive such universal, abstract representation as a fixed point of a RG transformation and we present numerical experiments corroborating this picture.
Abstraction has always [^18] According to both ChatGPT and Deepseek, as of October $19^{\rm th}$ 2025. been defined in relative terms, depending on what details are considered relevant. The process of abstraction connects different levels of detail in a hierarchy of representations. This process, we argue, can be described by an RG procedure that allows us to define absolute abstraction as the fixed point of the RG transformation, which describes the ultimate result of this process, in the limits of infinite detail or infinite generality. The distance from this limit allows us to define a quantitative measure of the level of abstraction of a representation as the Kullback-Leibler distance from the fixed point.
The fact that optimal representations of data rely on universal distributions is a well known fact in source coding. In fact, compression algorithms translate complex datasets into random strings of zeros and ones of maximal entropy [56]. This is also true of generative models such as Variational Auto-encoders [57] and Diffusion Models [58], that map the data to vectors of variables with a preassigned distribution. In these cases, as in source coding, information on the specific nature of the data is stored in the parameters that govern how the data is transformed. For example, in a neural network, the parameters of the different layers capture how each state of a deep representation is dressed by features at different scales, in order to generate data points. The internal representation – the code – just describes how the data is organized, and it is universal because it is the solution of an information theoretic optimization principle. In fact, the only common characteristic of data coming from very different domains is the coding cost, i.e. the number of bits needed to efficiently code each data point. The principle of maximal relevance [9] predicates that the coding cost should be as broadly distributed as possible, which in turn facilitates a robust alignment of different data sources along this dimension. The HFM arises as the ideal abstract representation because it satisfies the principle of maximal relevance, which qualifies it as an optimal scaffold for organizing data according to their coding cost [^19] Note that the coding cost $\log h_{n}(\mathbf{s})=gm_{\mathbf{s}}-\log Z_{n}$ depends linearly on the sufficient statistics $m_{\mathbf{s}}$, so the coding cost is itself a sufficient statistics..
@@
Understanding how the conceptual framework developed here can be extended to more complex domains such as language is an interesting avenue of further research [^21] In this respect, we note that the HFM can easily be generalised to variables $x_{i}$ taking value in an arbitrary set $\chi$ by invoking a transformation $\sigma:\chi\to\{0,1\}$ that maps each value of $x_{i}$ into a binary variable $s_{i}=\sigma(x_{i})$ [65]. Extending this analysis in the time dependent domain constitutes a considerably more challenging avenue.. In this paper we confine ourselves to the admittedly oversimplified domain of static representations of binary variables. Also, the range of breadth of the data that our numerical experiments explore is rather limited, as well as the representation capacity of the neural networks analysed. Extending this evidence further, and/or generalising the RG approach to more complex data structures, is an interesting avenue of further research.

### v5

@@ 5 Discussion
The main original contribution of this paper is to argue that, in layered information-processing systems such as artificial neural networks and the brain, universal representations emerge spontaneously from the combined effects of depth and breadth. We derive such universal, abstract representation as a fixed point of a RG transformation and we present numerical experiments corroborating this picture.
Abstraction is typically defined in relative terms, depending on what details are considered relevant. The process of abstraction connects different levels of detail in a hierarchy of representations. This process, we argue, can be described by an RG procedure that allows us to define absolute abstraction as the fixed point of the RG transformation, which describes the ultimate result of this process, in the limits of infinite detail or infinite generality. The distance from this limit allows us to define a quantitative measure of the level of abstraction of a representation as the Kullback-Leibler distance from the fixed point.
The fact that optimal representations of data rely on universal distributions is a well known fact in source coding. In fact, compression algorithms translate complex datasets into random strings of zeros and ones of maximal entropy [54]. This is also true of generative models such as Variational Auto-encoders [55] and Diffusion Models [56], that map the data to vectors of variables with a preassigned distribution. In these cases, as in source coding, information on the specific nature of the data is stored in the parameters that govern how the data is transformed. For example, in a neural network, the parameters of the different layers capture how each state of a deep representation is dressed by features at different scales, in order to generate data points. The internal representation – the code – just describes how the data is organized, and it is universal because it is the solution of an information theoretic optimization principle. In fact, the only common characteristic of data coming from very different domains is the coding cost, i.e. the number of bits needed to efficiently code each data point. The principle of maximal relevance [10] predicates that the coding cost should be as broadly distributed as possible, which in turn facilitates a robust alignment of different data sources along this dimension. The HFM arises as the ideal abstract representation because it satisfies the principle of maximal relevance, which qualifies it as an optimal scaffold for organizing data according to their coding cost [^18] Note that, in the HFM, the coding cost $\log h_{n}(\mathbf{s})=gm_{\mathbf{s}}-\log Z_{n}$ depends linearly on the sufficient statistics $m_{\mathbf{s}}$. So the coding cost is itself a sufficient statistics..
@@
Understanding how the conceptual framework developed here can be extended to more complex domains such as language is an interesting avenue of further research [^20] In this respect, we note that the HFM can easily be generalised to variables $x_{i}$ taking value in an arbitrary set $\chi$ by invoking a transformation $\sigma:\chi\to\{0,1\}$ that maps each value of $x_{i}$ into a binary variable $s_{i}=\sigma(x_{i})$ [63]. Extending this analysis in the time dependent domain constitutes a considerably more challenging avenue.. In this paper we confine ourselves to the admittedly oversimplified domain of static representations of binary variables. Also, the range of breadth of the data that our numerical experiments explore is rather limited, as well as the representation capacity of the neural networks analysed. Yet, these results align with empirical studies of large neural networks trained on heterogeneous datasets which exhibit a convergence toward shared representations [6]. This suggests that the RG approach discussed here may be extended to more complex data structures.